One number is 8 more than 1/2 of another number.The sum of the numbers is 23.What are the numbers?
Answer by Cromlix(4381)

Hi there,
Make 'another number' = x
'One number' = (8 + x/2)
Therefore, x + (8 + x/2) = 23
Remove brackets
x + 8 + x/2 = 23
Multiply throughout by 2
=> 2x + 16 + x = 46
Collect like terms
3x = 46 - 16
3x = 30
x = 10
Two numbers are:
10 and 13
Hope this helps:-)